OpenStudy (anonymous):

Use an inverse matrix to solve the equation or system: {x+2y=15} {2x+4y=30}

OpenStudy (anonymous):

Are we unable to find the inverse, or just not sure about any of it?

OpenStudy (anonymous):

I'm not sure how to find the inverse

OpenStudy (queelius):

First, to get a handle on this, you'll want to rewrite the system of equations into a matrix form. That is, Ax = b, where A is a 2x2 matrix, x is a vector [x, y], and b is a vector [15, 30]. Do you think you can give me what A is?

OpenStudy (queelius):

I should note that x and b are column vectors.

OpenStudy (anonymous):

Well, it's not hard for a 2x2 matrix fortunately. If you have a matrix \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\] the inverse is: \[\frac{ 1 }{ ab-bc }*\left[\begin{matrix}d & -b \\ -c & a\end{matrix}\right]\] Do you know how to put the equation given into a matrix equation?

OpenStudy (anonymous):

No I don't think so

OpenStudy (anonymous):

Well, the numbers that go into your matrix are the coefficients of x and y in each of your equations. So you have: x + 2y = 15 2x + 4y = 30. So the coefficients of the x's and y's will get plugged into the matrix, giving you: \[\left[\begin{matrix}1 & 2 \\ 2 & 4\end{matrix}\right]\]The question expects you to take this and find its inverse, which can be done using the formula from above.

OpenStudy (anonymous):

[4 -2] [2 1] ?

OpenStudy (anonymous):

\[\left[\begin{matrix}4 & -2 \\ -2 & 1\end{matrix}\right]\]So that's part of finding the inverse. I still need to divide that by ad-bc

OpenStudy (anonymous):

Once you do ad-bc and try to divide by that quantity, you should notice a big problem :)

OpenStudy (anonymous):

Is ad-bc 0?

OpenStudy (anonymous):

Yep. And since the formula requires you to divide by ad-bc, you'd be dividing by 0! That's a big no-no. What this tells you is that the matrix DOES NOT have an inverse.

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

Well, it still wanted you to get some sort of solution, right? The fact that this matrix doesn't have an inverse gives you two possibilities. I can either have no solution or infinitely many solutions. I think the question would still want you state which one of thse is the case

OpenStudy (anonymous):

nice one @Concentrationalizing